# Optimal Rates of Convergence for Entropy Regularization in Discounted Markov Decision Processes

### v3

@@ 2.2. State-action geometry of entropy regularization
which can be seen using the dominated convergence theorem [37]. Moreover, we have the classic characterization of the set of state-action distributions.
Proposition 2.4 (State-action polytope, [18]).
@@
Further, it is well known that for a state-action distribution $\nu\in\mathscrsfs{D}$ a policy $\pi\in\Delta_{\mathscr{A}}^{\mathscr{S}}$ with $\nu^{\pi}=\nu$ can be computed by conditioning, see for example [26, 39, 30], and hence we have
@@
Direct computation shows that $D_{A|S}$ is the Bregman divergence induced by $H_{A|S}$, see [42, A.1].
@@
This is a direct consequence of $R_{\tau}(\pi)=r^{\top}\nu^{\pi}-\tau D_{A|S}(\nu,{\nu_{0}})$, see also [42]. ∎
The Pythagorean theorem can be generalized to Bregman divergences, which is well known in the field of information geometry and convex optimization [16, 10, 4, 5]. As the Kakade divergence is not a Bregman divergence, it is not included in those general results, but using the characterization of $D_{\operatorname{K}}$ as the pullback of a conditional KL-divergence $D_{A|S}$ allows us to provide a Pythagorean theorem for the Kakade divergence and $s$-rectangular policy classes.
@@
Consider the set $D\coloneqq\{\nu^{\pi}:\pi\in\Pi\}\subseteq\mathscrsfs{D}$ of state-action distributions arising from the policy class $\Pi$, which is again a polytope [39, Remark 55]. In particular, $D$ is convex and we can pass to the corresponding state-action distributions $\nu_{0},\hat{\nu},\nu$ and apply the Pythagorean theorem for Bregman divergences, see Section A.2. If further $\Pi_{s}=\Delta_{\mathscr{A}}\cap\mathcal{L}_{s}$, then we have $D=\mathscrsfs{D}\cap\mathcal{L}$ for an affine subspace $\mathcal{L}$ [39, Proposition 14]. Consequently, the Bregman projection $\hat{\nu}$ will always lie at the relative interior $\operatorname{int}(D)$ and we get equality in the Pythagorean theorem. ∎
@@ 2.3. Regularized linear programs and Hessian gradient flows
We have seen that the state-action distributions of the optimal regularized policies $\pi^{\star}_{\tau}$ solve the linear program associated with the Markov decision process with a conditional entropic regularization. Here, we see that the solutions of regularized linear programs solve the gradient flow of the unregularized linear objective with respect to the Riemannian metric induced by the convex regularizer. We refer to [3] for a more general discussion of Hessian gradient flows, where slightly stronger assumptions on $\phi$ are made. We work in the following setting.
@@
Surely, we have $\hat{x}\in F^{\star}$ as $c^{\top}\hat{x}=\lim_{n\to\infty}c^{\top}x_{t_{n}}=\max_{x\in P}c^{\top}x$ by Corollary 2.10. Further, by the central path property we have for any optimizer $x^{\prime}\in F^{\star}$ that
@@ 3. Geometry and Sublinear Convergence of Kakade Gradient Flows
Our goal is to study the solutions $\pi_{\tau}^{\star}$ of the entropy-regularized reward $R_{\tau}$. For linear programs, we have seen in Section 2.3 that the solutions of the regularized problems solve the corresponding Hessian gradient flow. Recall, that the entropy-regularized reward optimization problem is equivalent to a linear program in state-action space with a conditional KL regularization. Hence, the state-action distributions solve a Hessian gradient flow and consequently, the optimal regularized policies solve a gradient flow with respect to some metric. In this section, we study this Riemannian metric on the space of policies and provide an explicit expression of the gradient dynamics.
@@ 3.1. The Kakade metric and policy gradient theorems
The following metric on the policy space was proposed in the context of natural gradients [25].
@@
The Kakade metric has been referred to as the Fisher-Rao metric on $\Delta_{\mathscr{A}}^{\mathscr{S}}$ as this reduces to the Fisher-Rao metric if $\lvert\mathscr{S}\rvert=1$, see [27]. We choose the name Kakade metric here, as there exist multiple extensions of the Fisher-Rao metric to products of simplicies. For example in a game-theoretic context, when considering independently chosen strategies of the players, it might be more natural to work with the product metric, meaning, the sum of the Fisher-Rao metrics over the individual factors, which corresponds to the pullback of the Fisher-Rao metric on the simplex of joint distributions under the independence model [36, 7]. Other weightings of the Fisher-Rao metrics over the individual factors are also possible, see [36] for an in-depth discussion of different choices and their invariance properties. Further, this specific Riemannian metric has been described as the limit of weighted Fisher-Rao metrics on the finite-horizon path spaces [6, 44, 51]. Note that although the Kakade metric is closely connected to the weighted entropy $H_{\operatorname{K}}$, it is not the Hessian metric of $H_{\operatorname{K}}$, in fact, it is not a Hessian metric at all [40, Remark 13].
In general, the Kakade metric is only a pseudo-metric and the following assumption ensures positive definiteness and we make it for the remainder of our analysis.
@@
Kakade gradient flows are well-posed, meaning that they admit a unique solution $(\pi_{t})_{t\in\mathbb{R}_{\geq 0}}$, both in the unregularized case [3, 40] and the regularized case [37, 27].
@@
and connects the gradient of the reward to the advantage function [48, 1]. Inspired by this, we provide an explicit formula for the gradient of the reward with respect to the Kakade metric.
@@
Lemma 3.4 (Derivatives of state-action distributions, [39]).
@@
Note that the gradient of the reward with respect to the Kakade metric is independent of the initial distribution $\mu\in\Delta_{\mathscr{S}}$ both for the regularized and the unregularized reward.
@@ 3.2. State-action geometry of Kakade gradient flows
Our goal is to use tools from Hessian gradient flows, but the Kakade metric is not a Hessian metric. However, it was shown that the policy polytope $\Delta_{\mathscr{A}}^{\mathscr{S}}$ endowed with the Kakade metric is isometric to the state-action polytope endowed with the Hessian metric induced by the conditional entropy [40]. This allows us to borrow from the results on Hessian gradient flows.
@@
It is elementary to check the convexity of $H_{A|S}$ on $\mathbb{R}_{>0}^{\mathscr{A}\times\mathscr{S}}$, but it also easily seen that $\nabla^{2}H_{A|S}$ has zero eigenvalues. However, it can be shown that $\nabla^{2}H_{A|S}$ is strictly definite on $T\mathscrsfs{D}$ and therefore induces a Riemannian metric on $\operatorname{int}(\mathscrsfs{D})$, see [40].
@@
Theorem 3.7 (State-action geometry of the Kakade metric, [40]).
@@
between policies and state-action distributions is an isometric diffeomorphism.
@@
Theorem 3.7 ensures that $\pi\mapsto\nu^{\pi}$ is an isometric diffeomorphism. Hence, by Lemma 3.8 we have that $(\pi_{t})_{t\geq 0}$ solves the Kakade gradient, if and only if $(\nu^{\pi_{t}})_{t\geq 0}$ solves the Hessian gradient flow with respect to the conditional Fisher-Rao metric. Hence, we can apply Section 2.3 and therefore obtain that both (ii) and (iii) hold on the maximal existence interval $[0,T_{\infty})$ of the gradient flow. Further, (iv) follows from Corollary 2.11 if $T_{\infty}=+\infty$, hence, it remains to show well-posedness.
The well-posedness of the conditional Fisher-Rao gradient flow of linear programs has been established in [37], which can be carried over to the Kakade gradient flow using the isometry. Here, we offer a different proof that utilizes the central path property of the Kakade gradient flow. The local well-posedness of the gradient flow follows from the Picard-Lindelöf theorem and it remains to show that $T_{\infty}\coloneqq\inf\{t>0:\inf_{s\in[0,t)}\operatorname{dist}(\pi_{s},\partial\Delta_{\mathscr{A}}^{\mathscr{S}})=0\}=+\infty$. Assume that $T_{\infty}<+\infty$, then it is elementary to check that $\pi_{t}\to\pi_{T_{\infty}}$ for $t\nearrow T_{\infty}$, where $\pi_{T_{\infty}}$ solves the entropy-regularized problem with strength $\tau=T_{\infty}^{-1}$. Note that $\pi_{T_{\infty}}\in\operatorname{int}(\Delta_{\mathscr{A}}^{\mathscr{S}})$ contradicting $\inf_{t\in[0,T_{\infty}]}\operatorname{dist}(\pi_{t},\partial\Delta_{\mathscr{A}}^{\mathscr{S}})=0$. ∎
@@
In the case of multiple optimal policies, the Kakade gradient flow will converge towards the Kakade projection of the initial policy to the set of optimal policies. The selection of gradient schemes of a particular optimizer in the case of multiple optima is commonly referred to as the implicit or algorithmic bias of the method. In the context of reinforcement learning the implicit bias was studied for discrete time natural policy gradient and policy mirror descent methods. First, for a natural actor-critic scheme in linear Markov decision processes the Kakade divergence of the optimization trajectory $(\pi_{k})_{k\in\mathbb{N}}$ to the maximum entropy policy $\pi^{\star}$ is bounded by $D_{\operatorname{K}}(\pi^{\star},\pi_{k})\leq\log k+(1-\gamma)^{-2}$ [21]. This control on the Bregman divergence to the maximum entropy policy ensures that the probability of selecting an optimal action $a\in A^{\star}_{s}$ decays at most like $\pi_{k}(a|s)\geq ck^{-1}$, but fails to identify the limiting policy.
On the other hand, a policy mirror descent scheme with decaying entropy regularization strength $\tau_{k}$ was studied in [32]. Here, it is shown that for several choices of regularization strengths and stepsizes the resulting policies $\pi_{k}$ converge to the maximum entropy optimal policy thereby characterizing the algorithmic bias of this approach. Whereas this analysis can characterize the limit of the optimization scheme it utilizes decaying explicit regularization.
@@
In other contexts, in particular for regression tasks, a huge variety of implicit bias results has been established for gradient and mirror descent [46, 20, 13, 8, 23, 15, 14]. In particular, in [22] the dynamics of gradient descent are shown to stay close to the solutions of norm-constrained optimization problems for a linear regression problem. This allows them to show convergence in direction of the gradient descent iterates towards the maximum-margin solution. In contrast, we study the reward optimization problem and natural policy gradient flows, which allows us to obtain a precise characterization of the optimization trajectory via the central path property.
@@
Let us fix $s\in\mathscr{S}$. By Theorem 3.3 the Kakade gradient flow $(\pi_{t})_{t\geq 0}$ is independent of the initial distribution $\mu$, as long as it has full support. Let us pick some $\mu_{n}$ with full support and $\mu_{n}\to\delta_{s}$, then $R^{\mu_{n}}(\pi)\to V^{\pi}(s)$ for any $\pi$. By Corollary 2.10 we have
@@ 4. Optimal Exponential Convergence Rates for Entropy Regularization
We now improve the sublinear convergence guarantee from Proposition 2.3 on the entropy regularization error. To this end we work with the interpretation of the solutions of the regularized problems as solutions of the Kakade gradient flow and employ the explicit expression (3.6) and the sublinear convergence guarantee from Proposition 3.11 to establish linear convergence of Kakade gradient flows. We complement this with a lower bound that matches the upper bound up to a polynomial factor. We work in the following setting.
@@ 4.1. Convergence in value
The above result ensures that the probability $\pi_{t}(a|s)$ of selecting a suboptimal action $a$ decays exponentially which implies exponential convergence of the reward $R(\pi_{t})$ achieved by the policies.
@@
Consider a discrete discounted Markov decision process. Then for any policy $\pi\in\Delta_{\mathscr{A}}^{\mathscr{S}}$ it holds that

### v5

@@ 2.2. State-action geometry of entropy regularization
which can be seen using the dominated convergence theorem [mueller2023thesis]. Moreover, we have the classic characterization of the set of state-action distributions.
Proposition 2.4 (State-action polytope, [derman1970finite]).
@@
Further, it is well known that for a state-action distribution $\nu\in\mathscrsfs{D}$, a policy $\pi\in\Delta_{\mathscr{A}}^{\mathscr{S}}$ with $\nu^{\pi}=\nu$ can be computed by conditioning, see for example [kallenberg1994survey, muller2022pomdps, laroche2023occupancy], and hence we have
@@
Direct computation shows that $D_{A|S}$ is the Bregman divergence induced by $H_{A|S}$, see [neu2017unified, A.1].
@@
This is a direct consequence of $R_{\tau}(\pi)=r^{\top}\nu^{\pi}-\tau D_{A|S}(\nu,{\nu_{0}})$, see also [neu2017unified]. ∎
The Pythagorean theorem can be generalized to Bregman divergences, which is well known in the field of information geometry and convex optimization [Csiszár2012, bubeck2015convex, amari2016information, ay2017information]. As the Kakade divergence is not a Bregman divergence, it is not included in those general results, but using the characterization of $D_{\operatorname{K}}$ as the pullback of a conditional KL-divergence $D_{A|S}$ allows us to provide a Pythagorean theorem for the Kakade divergence and $s$-rectangular policy classes.
@@
Consider the set $D\coloneqq\{\nu^{\pi}:\pi\in\Pi\}\subseteq\mathscrsfs{D}$ of state-action distributions arising from the policy class $\Pi$, which is again a polytope [muller2022pomdps, Remark 55]. In particular, $D$ is convex and we can pass to the corresponding state-action distributions $\nu_{0},\hat{\nu},\nu$ and apply the Pythagorean theorem for Bregman divergences, see Section A.2. If further $\Pi_{s}=\Delta_{\mathscr{A}}\cap\mathcal{L}_{s}$, then we have $D=\mathscrsfs{D}\cap\mathcal{L}$ for an affine subspace $\mathcal{L}$ [muller2022pomdps, Proposition 14]. Consequently, the Bregman projection $\hat{\nu}$ will always lie at the relative interior $\operatorname{int}(D)$ and we get equality in the Pythagorean theorem. ∎
@@ 2.3. Regularized linear programs and Hessian gradient flows
We have seen that the state-action distributions of the optimal regularized policies $\pi^{\star}_{\tau}$ solve the linear program associated with the Markov decision process with a conditional entropic regularization. Here, we see that the solutions of regularized linear programs solve the gradient flow of the unregularized linear objective with respect to the Riemannian metric induced by the convex regularizer. We refer to [alvarez2004hessian] for a more general discussion of Hessian gradient flows, where slightly stronger assumptions on $\phi$ are made. We work in the following setting.
@@
Surely, we have $\hat{x}\in F^{\star}$ as $c^{\top}\hat{x}=\lim_{n\to\infty}c^{\top}x_{t_{n}}=\max_{x\in P}c^{\top}x$ by Corollary 2.10. Further, by the central path property, we have for any optimizer $x^{\prime}\in F^{\star}$ that
@@ 3. Geometry and Sublinear Convergence of Kakade Gradient Flows
Our goal is to study the solutions $\pi_{\tau}^{\star}$ of the entropy-regularized reward $R_{\tau}$. For linear programs, we have seen in Section 2.3 that the solutions of the regularized problems solve the corresponding Hessian gradient flow. Recall that the entropy-regularized reward optimization problem is equivalent to a linear program in state-action space with a conditional KL regularization. Hence, the state-action distributions solve a Hessian gradient flow and consequently, the optimal regularized policies solve a gradient flow with respect to some metric. In this section, we study this Riemannian metric on the space of policies and provide an explicit expression of the gradient dynamics.
@@ 3.1. The Kakade metric and policy gradient theorems
The following metric on the policy space was proposed in the context of natural gradients [kakade2001natural].
@@
The Kakade metric has been referred to as the Fisher-Rao metric on $\Delta_{\mathscr{A}}^{\mathscr{S}}$ as this reduces to the Fisher-Rao metric if $\lvert\mathscr{S}\rvert=1$, see [kerimkulov2023fisher]. We choose the name Kakade metric here, as there exist multiple extensions of the Fisher-Rao metric to products of simplices. For example, in a game-theoretic context, when considering independently chosen strategies of the players, it might be more natural to work with the product metric, meaning the sum of the Fisher-Rao metrics over the individual factors, which corresponds to the pullback of the Fisher-Rao metric on the simplex of joint distributions under the independence model [montufar2014fisher, boll2024geometric]. Other weightings of the Fisher-Rao metrics over the individual factors are also possible, see [montufar2014fisher] for an in-depth discussion of different choices and their invariance properties. Further, this specific Riemannian metric has been described as the limit of weighted Fisher-Rao metrics on the finite-horizon path spaces [bagnell2003covariant, peters2003reinforcement, wolfer2023information]. Note that although the Kakade metric is closely connected to the weighted entropy $H_{\operatorname{K}}$, it is not the Hessian metric of $H_{\operatorname{K}}$, in fact, it is not a Hessian metric at all [muller2024geometry, Remark 13].
In general, the Kakade metric is only a pseudo-metric. The following assumption ensures positive definiteness, which we assume for the remainder of our analysis.
@@
Kakade gradient flows are well-posed, meaning that they admit a unique solution $(\pi_{t})_{t\in\mathbb{R}_{\geq 0}}$, both in the unregularized case [alvarez2004hessian, muller2024geometry] and the regularized case [mueller2023thesis, kerimkulov2023fisher].
@@
and connects the gradient of the reward to the advantage function [sutton1999policy, agarwal2021theory]. Inspired by this, we provide an explicit formula for the gradient of the reward with respect to the Kakade metric.
@@
Lemma 3.4 (Derivatives of state-action distributions, [muller2022pomdps]).
@@
Note that the gradient of the reward with respect to the Kakade metric is independent of the initial distribution $\mu\in\Delta_{\mathscr{S}}$, both for the regularized and the unregularized reward.
@@ 3.2. State-action geometry of Kakade gradient flows
Our goal is to use tools from Hessian gradient flows, but the Kakade metric is not a Hessian metric. However, it was shown that the policy polytope $\Delta_{\mathscr{A}}^{\mathscr{S}}$ endowed with the Kakade metric is isometric to the state-action polytope endowed with the Hessian metric induced by the conditional entropy [muller2024geometry]. This allows us to borrow from the results on Hessian gradient flows.
@@
It is elementary to check the convexity of $H_{A|S}$ on $\mathbb{R}_{>0}^{\mathscr{A}\times\mathscr{S}}$, but it also easily seen that $\nabla^{2}H_{A|S}$ has zero eigenvalues. However, it can be shown that $\nabla^{2}H_{A|S}$ is strictly definite on $T\mathscrsfs{D}$ and therefore induces a Riemannian metric on $\operatorname{int}(\mathscrsfs{D})$, see [muller2024geometry].
@@
Theorem 3.7 (State-action geometry of the Kakade metric, [muller2024geometry]).
@@
between policies and state-action distributions is an isometric diffeomorphism, where the conditioning map is defined in (2.22). On the whole conditional probability polytope, $\Xi\colon\Delta_{\mathscr{A}}^{\mathscr{S}}\to\mathscrsfs{D}$, $\pi\mapsto\nu^{\pi}$ is a continuous bijection with continuous inverse.
@@
We set
$\pi_{t}=\operatorname*{arg\,max}_{\pi\in\Delta_{\mathscr{A}}^{\mathscr{S}}}\Big\{R(\pi)-t^{-1}D_{\operatorname{K}}(\pi,\pi_{0})\Big\},$ (3.16)
where it is elementary to check $\pi_{t}\in\operatorname{int}(\Delta_{\mathscr{A}}^{\mathscr{S}})$. By Proposition 2.6, the element $\nu_{t}\coloneqq\nu^{\pi_{t}}$ solves the regularized linear program with regularization strength $t^{-1}$ and by Proposition 2.9 $(\nu_{t})_{t\geq 0}$ solves the Hessian gradient flow with respect to the conditional Fisher-Rao metric. As $\pi\mapsto\nu^{\pi}$ is an isometric diffeomorphism by Theorem 3.7, Lemma 3.8 ensures that $(\pi_{t})_{t\geq 0}$ solves the Kakade gradient flow. This shows both (i). and (ii).. For (iii)., we pick $\pi^{\star}\in\operatorname*{arg\,min}_{\pi\in\Pi^{\star}}D_{{\operatorname{K}}}(\pi^{\star},\pi_{0})$ and set $\nu^{\star}\coloneqq\nu^{\pi^{\star}}$. By Proposition 2.6 we have use Corollary 2.10 to obtain
$R^{\star}-R(\pi_{t})=r^{\top}\nu^{\star}-r^{\top}\nu_{t}\leq\frac{D_{A|S}(\nu^{\star},\nu_{0})-D_{A|S}(\nu_{t},\nu_{0})}{t}=\frac{D_{{\operatorname{K}}}(\pi^{\star},\pi_{0})-D_{{\operatorname{K}}}(\pi_{t},\pi_{0})}{t}.$
Finally, for (iv). we use Corollary 2.11 together with the continuity of the inverse of the state-action map $\Xi$ and obtain
$\lim_{t\to+\infty}\pi_{t}=\Xi^{-1}(\lim_{t\to+\infty}\nu_{t})=\Xi^{-1}(\nu^{\star})=\pi^{\star}.$
@@
In the case of multiple optimal policies, the Kakade gradient flow will converge towards the Kakade projection of the initial policy to the set of optimal policies. The selection of gradient schemes of a particular optimizer in the case of multiple optima is commonly referred to as the implicit or algorithmic bias of the method. In the context of reinforcement learning, the implicit bias was studied for discrete-time natural policy gradient and policy mirror descent methods. First, for a natural actor-critic scheme in linear Markov decision processes the Kakade divergence of the optimization trajectory $(\pi_{k})_{k\in\mathbb{N}}$ to the maximum entropy policy $\pi^{\star}$ is bounded by $D_{\operatorname{K}}(\pi^{\star},\pi_{k})\leq\log k+(1-\gamma)^{-2}$ [hu2021actor]. This control on the Bregman divergence to the maximum entropy policy ensures that the probability of selecting an optimal action $a\in A^{\star}_{s}$ decays at most like $\pi_{k}(a|s)\geq ck^{-1}$, but fails to identify the limiting policy.
On the other hand, a policy mirror descent scheme with decaying entropy regularization strength $\tau_{k}$ was studied in [li2023homotopic]. Here, it is shown that for several choices of regularization strengths and stepsizes, the resulting policies $\pi_{k}$ converge to the maximum entropy optimal policy, thereby characterizing the algorithmic bias of this approach. Whereas this analysis can characterize the limit of the optimization scheme, it utilizes decaying explicit regularization.
@@
In other contexts, in particular for regression tasks, a huge variety of implicit bias results have been established for gradient and mirror descent [soudry2018implicit, gunasekar2018characterizing, chizat2020implicit, bowman2022implicit, jin2023implicit, cohen2023deep, chou2024gradient]. In particular, in [pmlr-v125-ji20a], the dynamics of gradient descent are shown to stay close to the solutions of norm-constrained optimization problems for a linear regression problem. This allows them to show convergence in the direction of the gradient descent iterates towards the maximum-margin solution. In contrast, we study the reward optimization problem and natural policy gradient flows, which allows us to obtain a precise characterization of the optimization trajectory via the central path property.
@@
Let us fix $s\in\mathscr{S}$. By Theorem 3.3, the Kakade gradient flow $(\pi_{t})_{t\geq 0}$ is independent of the initial distribution $\mu$, as long as it has full support. Let us pick some $\mu_{n}$ with full support and $\mu_{n}\to\delta_{s}$, then $R^{\mu_{n}}(\pi)\to V^{\pi}(s)$ for any $\pi$. By Corollary 2.10 we have
@@ 4. Optimal Exponential Convergence Rates for Entropy Regularization
We now improve the sublinear convergence guarantee from Proposition 2.3 on the entropy regularization error. To this end, we work with the interpretation of the solutions of the regularized problems as solutions of the Kakade gradient flow and employ the explicit expression (3.6) and the sublinear convergence guarantee from Proposition 3.11 to establish linear convergence of Kakade gradient flows. We complement this with a lower bound that matches the upper bound up to a polynomial factor. We work in the following setting.
@@ 4.1. Convergence in value
The above result ensures that the probability $\pi_{t}(a|s)$ of selecting a suboptimal action $a$ decays exponentially, which implies exponential convergence of the reward $R(\pi_{t})$ achieved by the policies.
@@
Consider a discrete discounted Markov decision process. Then, for any policy $\pi\in\Delta_{\mathscr{A}}^{\mathscr{S}}$ it holds that